# Modelling Micro-Doppler Signature of Drone Propellers in Distributed ISAC

### v2

@@ I Introduction
In order to fill this gap and develop adequate techniques, it is important to have access to a large amount of target data. Bistatic Radar measurement facility (BiRa), installed at the Ilmenau University of Technology, can play an interesting role in the way to solve this problem, since it is able to automatically perform multiple multiaspect micro-Doppler measurements along a defined set of bistatic geometry positions, with different parameters, and without user intervention [1]. This allows collecting numerous measurement data in a single unmanned measurement run.
@@ II-A Classic Propeller Micro-Doppler Model
If a scatter point target P has a range $R_{P}^{(m)}(t)$ with respect to a monostatic radar, its returns to this radar can be expressed as [4]:
$y_{P}(t)=\gamma_{P}(t)\exp\bigg{\{}j\omega_{0}\left[t-\frac{2R_{P}^{(m)}(t)}{c}\right]\bigg{\}}+z(t)$ (2)
@@
$\gamma_{P}(t)=\gamma_{0}\sqrt{\frac{c\sigma_{P}}{(4\pi^{3})R_{P}(t)^{4}f_{0}^{2}}}$ (3)
with $\sigma_{P}$ being the RCS of the scatter point $P$.
Since the wavelengths involved in the proposed ISAC systems are smaller than the dimensions of most drone structures, except for some micro and nano drones which are out of the scope of this study, the electromagnetic scattering of these targets behave similarly to what happens in the optical area.
@@
$\begin{split}\tilde{h}_{i}(t)&=\sum_{i=1}^{N_{B}}\bigg{\{}\tilde{\gamma}_{i}(t)e^{\left[j\frac{\omega_{0}}{c}\left(-2R_{O}(t)+L_{B}\cos(\omega t+\varphi_{B}(i))\cos(\beta)\right)\right]}\cdot\\ &\cdot L\operatorname{sinc}\left[\frac{\omega_{0}L_{B}}{c}\cos(\omega t+\varphi_{B}(i))\cos(\beta)\right]\bigg{\}}+\tilde{z}(t)\end{split}$ (4)
@@
$\varphi_{B}(i)=\varphi_{0}+2\pi\frac{i-1}{N_{B}}$ (5)
@@ II-B Bistatic Geometry
In order to be able to derive a similar model for the broadband as well as bistatic case, it is first necessary to have a geometrical model of the bistatic range of a single rotating point.
@@
Such a math model was presented in [8], where the author develops the mathematical formulation of the bistatic range of a scatter point P, in a bistatic radar environment, rotating around a point O, according to Fig. 1. This bistatic range is given by
$R_{P}^{(b)}(t)\approx R_{O}-A_{B}l\cos(\omega t+\varphi_{B})$ (6)
where $R_{O}=R_{T}+R_{R}$ is the total bistatic range of the rotation center O, and
$A_{B}=\sqrt{4\cos(\beta/2)^{2}-(\cos\beta_{T}+\cos\beta_{R})^{2}}$ (7)
$\varphi_{B}=\varphi_{T}-\arctan\left(\frac{\sin\beta_{R}\sin(\varphi_{T}-\varphi_{R})}{\sin\beta_{T}+\sin\beta_{R}\cos(\varphi_{T}-\varphi_{R})}\right)$ (8)
Based on (6), and using the same method as in Section II-A for the mono-static case, it is now possible to derive a model for bistatic returns of a rotor with $N_{B}$ rotating blades of length $L_{B}$ each.
@@ II-C OFDM Waveform
The model from equation (1) assumes a simple CW waveform with center frequency $f_{0}$, where $\omega_{0}=2\pi f_{0}$. This formulation can be considered a good approximation for radars with narrow bandwidth. However, for broadband systems this formulation might not properly reproduce the signal features, since the same target scatter point can produce different Doppler contributions due to the different frequency components of the transmitted signal.
@@ II-D Bistatic OFDM Mathematical Formulation
$x(\mu,m)=\sum_{n=1}^{N}D(n,m)\exp\left(j\omega_{n}(m+\frac{\mu}{N})T\right)$ (9)
@@
Additionally, ${D(n,m)}$ is the complex amplitude for each subcarrier, also called complex modulation symbol, given by the modulation technique, e.g. phase-shift keying (PSK).
@@
So, as $\exp(j2\pi nm)=1$, and using $t$ in the expression as a matter of simplicity, (9) can be rewritten as
$x(\mu,m)=\exp(j\omega_{0}t)\sum_{n=1}^{N}D(n,m)\exp(j2\pi\frac{n\mu}{N})$ (11)
@@
$\begin{split}y_{P}^{(b)}(t,\tau)=&\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\bigg{[}j\omega_{n}\left(t-\frac{R_{P}^{(b)}(t)}{c}\right)\bigg{]}.\\ &.rect\left(\frac{\tau-\frac{R_{P}^{(b)}(t)}{c}}{T}-\frac{1}{2}\right)\Bigg{\}}+z(t)\end{split}$ (12)
@@
Now, if $P$ is a rotating scatter point, the bistatic range $R_{P}^{(b)}(t)$ is given by (6).
Once again, considering a blade $i$ as a line of infinite point scatters, each of its returns can be calculated, by using
$\begin{split}&y_{i}^{(b)}(t,\tau)=\int_{0}^{L_{B}}\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\bigg{[}j\omega_{n}\cdot\\ &\cdot\left(t-\frac{R_{P}^{(b)}(t)}{c}\right)\bigg{]}\operatorname{rect}\left(\frac{\tau-\frac{R_{P}^{(b)}(t)}{c}}{T}-\frac{1}{2}\right)\Bigg{\}}\,dl+z(t)\end{split}$ (13)
@@
$\begin{split}&y_{i}^{(b)}(t,\tau)=\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\left(j\omega_{n}t\right)\cdot\\ &\cdot\int_{l_{1}}^{l_{2}}{\exp\bigg{[}-j\frac{\omega_{n}}{c}(R_{P}^{(b)}(t))\bigg{]}}{dl}\Bigg{\}}+z(t)\end{split}$ (14)
@@
$\begin{split}&y^{(b)}(\mu,m)=\sum_{n=1}^{N}\Bigg{\{}D(n,m)\exp\left(j2\pi\frac{n\mu}{N}\right)\cdot\\ &\cdot\sum_{i=1}^{N_{B}}\biggl{[}\gamma_{ni}(m)e^{j\frac{\omega_{n}}{c}\left(-R_{O}(t)+\Delta R_{+}(\mu,i)\right)}\cdot\\ &\cdot\frac{l_{2}(\mu)-l_{1}(\mu)}{2}\operatorname{sinc}\left(\frac{\omega_{n}}{c}\Delta R_{-}(\mu,i)\right)\biggr{]}\Bigg{\}}+z(t)\end{split}$ (16)
@@
$\Delta R_{+}(\mu,i)=A_{B}\frac{l_{2}(\mu)+l_{1}(\mu)}{2}\cos(\omega t+\varphi_{B}(i))$ (17)
$\Delta R_{-}(\mu,i)=A_{B}\frac{l_{2}(\mu)-l_{1}(\mu)}{2}\cos(\omega t+\varphi_{B}(i))$ (18)
@@
$\bigg{\{}0,\frac{R_{O}(t)-(\frac{\mu}{N}-1)cT}{A_{B}\cos(\omega t+\varphi_{B}(i))},\frac{R_{O}(t)-\frac{\mu}{N}cT}{A_{B}\cos(\omega t+\varphi_{B}(i))}\bigg{\}}$ (19)
$\bigg{\{}L_{B},\frac{R_{O}(t)-(\frac{\mu}{N}-1)cT}{A_{B}\cos(\omega t+\varphi_{B}(i))},\frac{R_{O}(t)-\frac{\mu}{N}cT}{A_{B}\cos(\omega t+\varphi_{B}(i))}\bigg{\}}$ (20)
@@ III-B Bistatic Micro-Doppler Measurement Setup
For micro-Doppler analysis, the signal was processed in slow-time, i.e., a vector of the returns of the target in a single range bin, along different symbols. Some subsampling can also be applied in order to reduce the amount of data and the slow-time sampling frequency, and thus, to achieve a better visualization of the data, focusing the analysis on the bandwidth of interest. In this work, a subsampling factor of 8 means that we use only every eighth symbol available.
@@ III-C Measurements Analysis
We can also see that this sequence of impulses ends in the Doppler frequencies equivalent to the target’s maximum positive and negative relative speeds. In the current measurement, the maximum velocity is given by $v=2\pi f_{rot}l=2\pi\times 25\times 0.1655=641.25$ m/s, where $l=0.1655$ m is the radius of the propeller. Bistatic Doppler is given by [11]:
@@
where $\beta$ is the bistatic angle and $\delta$ the angle between the bistatic bisector, and $v$ the direction of target motion. In this case, we have $\delta$ = 0 and $\beta$ = 60°. So ${f_{D}}_{max}$= 555.4 Hz.
@@ IV-A Low Resolution Case
Fig. 5 presents a comparison between two micro-Doppler signatures: one simulated using the model developed in Section II, and one obtained from the measurement described in Section III. Both use the parameters described in Table I Setup 1, and geometrical configuration with bistatic angles of 30°.
@@ IV-B High Range Resolution Case
It is also possible to simulate a drone with multiple propellers by setting the coordinates of the rotation center point O (Fig. 1) of each rotor according to the drone geometry, generating each propeller returns using the proposed model, and then summing the returns from each propeller and from the static parts of the drone. Propellers and static parts return intensity can be adjusted by changing the RCS parameter $\sigma_{P}$ in (3). Future work will present a more detailed analysis of multi-propeller drone simulation using this model.
Fig. 8 shows a comparison between the range-Doppler signatures of a drone with two propellers from two sources: simulated from the proposed model and retrieved from BiRa measurements. Both cases use the parameters from Table I Setup 2, thus represent a high range resolution scenario. The main properties of the drone micro-Doppler signatures, such as range of the propellers, Doppler spread, and distance between impulses in frequency direction are the same on both sides.
@@ IV-C Advantages and Disadvantages
Many other methods can be used for modeling drone propellers, such as Finite Element Method (FEM), ray tracing, primitive-based modeling from finite discrete scatter points, and classic primitive continuous modeling. Compared to other methods, the most important advantages of the proposed method are its simplicity, low computational cost, and ability to represent the HRR case, while taking the waveform into account. Its main disadvantage is that it doesn’t take into account some electromagnetic phenomena, such as diffraction and multipath, which can introduce inaccuracies. However, its simplicity makes it suitable for applications where a large amount of simulated data with various parameters is required.
@@ V Conclusions and Outlook
Future work can comprise the use of simulated data using this model to train target classification algorithms, and the validation of this training method using measurement data.
@@ Acknowledgment
The authors would like to thank Dr.Ing. Tobias Nowack and M.Sc. Masoumeh Pourjafarian for their support throughout the BiRa measurement in ThiMo. This research is funded by the BMBF project 6G-ICAS4Mobility, Project No.16KISK241, by the Federal State of Thuringia, Germany, and by the European Social Fund (ESF) under grants 2017 FGI 0007 (project "BiRa") and 2021 FGI 0007 (project "Kreatör").

### v3

@@ I Introduction
In order to fill this gap and develop adequate techniques, it is important to have access to a large amount of target data. Bistatische-Radar-Messeinrichtung (BiRa) measurement facility, installed at the Ilmenau University of Technology, can play an interesting role in the way to solve this problem, since it is able to automatically perform multiple multiaspect micro-Doppler measurements along a defined set of bistatic geometry positions, with different parameters, and without user intervention [1]. This allows collecting numerous measurement data in a single unmanned measurement run.
@@ II-A Classic Propeller Micro-Doppler Model
If a scatter point target P has a range $R_{P}^{(m)}(t)$ with respect to a monostatic radar, its returns to this radar can be expressed as [4]
$y_{P}(t)=\gamma_{P}(t)\exp\bigg{\{}j\omega_{0}\left[t-\frac{2R_{P}^{(m)}(t)}{c}\right]\bigg{\}}+z(t),$ (2)
@@
$\gamma_{P}(t)=\gamma_{0}\sqrt{\frac{c\sigma_{P}}{(4\pi^{3})R_{P}(t)^{4}f_{0}^{2}}},$ (3)
$\sigma_{P}$ is the radar cross section (RCS) of the scatter point $P$.
Since the wavelengths in the proposed ISAC systems (Table I) are smaller than the dimensions of most drone structures, except for some micro and nano drones which are out of the scope of this study, the electromagnetic scattering of these targets behave similarly to what happens in the optical area.
@@
$\begin{split}\tilde{h}_{i}(t)&=\sum_{i=1}^{N_{B}}\bigg{\{}\tilde{\gamma}_{i}(t)e^{\left[j\frac{\omega_{0}}{c}\left(-2R_{O}(t)+L_{B}\cos(\omega t+\varphi_{B}(i))\cos(\beta)\right)\right]}\\ &\cdot L\operatorname{sinc}\left[\frac{\omega_{0}L_{B}}{c}\cos(\omega t+\varphi_{B}(i))\cos(\beta)\right]\bigg{\}}+\tilde{z}(t),\end{split}$ (4)
@@
$\varphi_{B}(i)=\varphi_{0}+2\pi\frac{i-1}{N_{B}}.$ (5)
@@ II-B Bistatic Geometry
In order to be able to derive a similar model for the broadband bistatic case, it is first necessary to have a geometrical model of the bistatic range of a single rotating point.
@@
Such a math model was presented in [8], where the author develops the mathematical formulation of the bistatic range of a scatter point P, in a bistatic radar environment, rotating around a point O, according to Fig. 1. This bistatic range is
$R_{P}^{(b)}(t)\approx R_{O}-A_{B}l\cos(\omega t+\varphi_{B}),$ (6)
where $R_{O}=R_{T}+R_{R}$ is the total bistatic range of the rotation center O, $A_{B}=\sqrt{4\cos(\beta/2)^{2}-(\cos\beta_{T}+\cos\beta_{R})^{2}}$, and
$\varphi_{B}=\varphi_{T}-\arctan\left(\frac{\sin\beta_{R}\sin(\varphi_{T}-\varphi_{R})}{\sin\beta_{T}+\sin\beta_{R}\cos(\varphi_{T}-\varphi_{R})}\right).$ (7)
Based on (6), and using the same method as in Section II-A for the monostatic case, it is now possible to derive a model for bistatic returns of a rotor with $N_{B}$ rotating blades of length $L_{B}$ each.
@@ II-C OFDM Waveform
The model from (1) assumes a simple CW waveform with center frequency $f_{0}$, where $\omega_{0}=2\pi f_{0}$. This formulation can be considered a good approximation for radars with narrow bandwidth. However, for broadband systems this formulation might not properly reproduce the signal features, since the same target scatter point can produce different Doppler contributions due to the different frequency components of the transmitted signal.
@@ II-D Bistatic OFDM Mathematical Formulation
$x(\mu,m)=\sum_{n=1}^{N}D(n,m)\exp\left(j\omega_{n}(m+\frac{\mu}{N})T\right),$ (8)
@@
Additionally, D(n,m){D(n,m)} is the complex amplitude for each subcarrier, also called complex modulation symbol, given by the modulation technique, e.g. phase-shift keying (PSK).
@@
So, as $\exp(j2\pi nm)=1$, and using $t$ in the expression as a matter of simplicity, ((8)) can be rewritten as
$x(\mu,m)=\exp(j\omega_{0}t)\sum_{n=1}^{N}D(n,m)\exp(j2\pi\frac{n\mu}{N}).$ (10)
@@
$y_{P}^{(b)}(t,\tau)=$ $\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\bigg{[}j\omega_{n}\left(t-\frac{R_{P}^{(b)}(t)}{c}\right)\bigg{]}$
$\cdot\operatorname{rect}\left(\frac{\tau-\frac{R_{P}^{(b)}(t)}{c}}{T}-\frac{1}{2}\right)\Bigg{\}}+z(t).$ (11)
@@
Now, if $P$ is a rotating scatter point, the bistatic range $R_{P}^{(b)}(t)$ is given by ((6)).
Once again, considering a blade $i$ as a line of infinite point scatters, each of its returns can be calculated by using
$y_{i}^{(b)}(t,\tau)=$ $\int_{0}^{L_{B}}\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\bigg{[}j\omega_{n}\left(t-\frac{R_{P}^{(b)}(t)}{c}\right)\bigg{]}$
$\cdot\operatorname{rect}\left(\frac{\tau-\frac{R_{P}^{(b)}(t)}{c}}{T}-\frac{1}{2}\right)\Bigg{\}}\,dl+z(t).$ (12)
@@
$\begin{split}y_{i}^{(b)}(t,\tau)=&\sum_{n=1}^{N}\Bigg{\{}\gamma_{n}D(n,m)\exp\left(j\omega_{n}t\right)\\ &\cdot\int_{l_{1}}^{l_{2}}{\exp\bigg{[}-j\frac{\omega_{n}}{c}(R_{P}^{(b)}(t))\bigg{]}}{dl}\Bigg{\}}+z(t).\end{split}$ (13)
@@
$\begin{split}&y^{(b)}(\mu,m)=\sum_{n=1}^{N}\Bigg{\{}D(n,m)\exp\left(j2\pi\frac{n\mu}{N}\right)\\ &\cdot\sum_{i=1}^{N_{B}}\biggl{[}\gamma_{ni}(m)e^{j\frac{\omega_{n}}{c}\left(-R_{O}(t)+\Delta R_{+}(\mu,i)\right)}\\ &\cdot\frac{l_{2}(\mu)-l_{1}(\mu)}{2}\operatorname{sinc}\left(\frac{\omega_{n}}{c}\Delta R_{-}(\mu,i)\right)\biggr{]}\Bigg{\}}+z(t),\end{split}$ (15)
@@
$\Delta R_{+}(\mu,i)=A_{B}\frac{l_{2}(\mu)+l_{1}(\mu)}{2}\cos(\omega t+\varphi_{B}(i)),$ (16)
$\Delta R_{-}(\mu,i)=A_{B}\frac{l_{2}(\mu)-l_{1}(\mu)}{2}\cos(\omega t+\varphi_{B}(i)),$ (17)
@@
$\bigg{\{}0,\frac{R_{O}(t)-(\frac{\mu}{N}-1)cT}{A_{B}\cos(\omega t+\varphi_{B}(i))},\frac{R_{O}(t)-\frac{\mu}{N}cT}{A_{B}\cos(\omega t+\varphi_{B}(i))}\bigg{\}},$ (18)
$\bigg{\{}L_{B},\frac{R_{O}(t)-(\frac{\mu}{N}-1)cT}{A_{B}\cos(\omega t+\varphi_{B}(i))},\frac{R_{O}(t)-\frac{\mu}{N}cT}{A_{B}\cos(\omega t+\varphi_{B}(i))}\bigg{\}}.$ (19)
@@ III-B Bistatic Micro-Doppler Measurement Setup
For micro-Doppler analysis, an OFDM radar processing was performed, as in [5], with a back to back measurement as reference for proper measurement calibration. The signal was then processed in slow-time, i.e., a vector of the returns of the target in a single range bin, along different symbols.
Some subsampling can also be applied in order to reduce the amount of data, while focusing the analysis on the bandwidth of interest. In this work, a subsampling factor of 8 means that we use only every eighth symbol available.
@@ III-C Measurements Analysis
We can also see that this sequence of impulses ends in the Doppler frequencies equivalent to the target’s maximum positive and negative relative speeds. In the current measurement, the maximum velocity is given by $v=2\pi f_{rot}l=2\pi\times 25\times 0.1655=26$ m/s, where $l=0.1655$ m is the radius of the propeller. Bistatic Doppler is given by [11]:
@@
where $\beta$ is the bistatic angle and $\delta$ the angle between the bistatic bisector, and $v$ the direction of target motion. In this case, we have $\delta$ = 0° and $\beta$ = 60°. So ${f_{D}}_{max}$= 555.4 Hz.
@@ IV-A Low Resolution Case
Fig. 5 presents a comparison between two micro-Doppler signatures: one simulated using the model developed in Section II, and one obtained from the measurement described in Section III. Both use the parameters described in Table I Setup 1, and geometrical configuration with bistatic angles of 60°.
@@ IV-B High Range Resolution Case
It is also possible to simulate a drone with multiple propellers by setting the coordinates of the rotation center point O (Fig. 1) of each rotor according to the drone geometry, generating each propeller returns using the proposed model, and then summing the returns from each propeller and from the static parts of the drone.
Fig. 8 shows a comparison between the range-Doppler signatures of a drone with two propellers from two sources: simulated from the proposed model and retrieved from BiRa measurements. Both cases use the parameters from Table I Setup 2, thus represent a high range resolution scenario. The main properties of the drone micro-Doppler signatures for the sake of classification, such as number of propellers, position and Doppler spread of each propeller, and distance between impulses in frequency direction are the same on both sides. In order to improve intensity match between simulation and measurement, a better RCS model for propellers and static parts returns can be included, by adjusting $\sigma_{P}$ in (3).
@@ IV-C Advantages and Disadvantages
Many other methods can be used for modeling drone propellers, such as Finite Element Method (FEM), ray tracing, primitive-based modeling from finite discrete scatter points, and classic primitive continuous modeling. Compared to other methods, the most important advantages of the proposed method are its simplicity, low computational cost, and ability to represent the HRR case, while taking the waveform into account. Among the disadvantages, it does not take into account some electromagnetic phenomena, such as diffraction and multipath, and it also cannot reproduce more complex phenomena such as frequency components due to vibration. However, its simplicity makes it suitable for applications where a large amount of simulated data with various parameters is required.
@@ V Conclusions and Outlook
Future work can comprise the addition of a more realistic reflectivity model for the main body of the target in the HRR case, the use of simulated data using this model to train target classification algorithms, and the validation of this training method using measurement data.
@@ Acknowledgment
The authors would like to thank Dr.Ing. Tobias Nowack and M.Sc. Masoumeh Pourjafarian for their support throughout the BiRa measurement in ThiMo. This research is funded by the BMBF project 6G-ICAS4Mobility, Project No.16KISK241, by the Federal State of Thuringia, Germany, and by the European Social Fund (ESF) under grants 2017 FGI 0007 (project ”BiRa”) and 2021 FGI 0007 (project ”Kreatör”).